OpenStudy (anonymous):

if log3(4x +1) =2 what is the value of x?

OpenStudy (anonymous):

\[\log _{3} ( 4x+1) = 2\]

OpenStudy (loser66):

any idea?

OpenStudy (anonymous):

no

OpenStudy (loser66):

neither do I

OpenStudy (anonymous):

rewrite in equivalent exponential form \[\log_b(A)=C \iff b^C=A\] then solve for \(x\)

OpenStudy (loser66):

thank you, now, I know

OpenStudy (souvik):

now 4x+1=3^2...now you can do it from here...

OpenStudy (anonymous):

i got 2 as my answer.

OpenStudy (souvik):

right...

OpenStudy (loser66):

yup

OpenStudy (anonymous):

thank you very much !